OpenStudy (anonymous):

calculate ! sqrt(1+97x98x99100) by using algebraic concept

OpenStudy (anonymous):

is it sqrt(1+97x98x99x100)? or is the last number really 99100?

OpenStudy (anonymous):

opppss, sorry you are right, sqrt(1+97x98x99x100)

OpenStudy (anonymous):

alright, hold on...

OpenStudy (anonymous):

Let x=101 \[\sqrt{1+(x-4)(x-3)(x-2)(x-1)}\] Rearranging, \[=\sqrt{(x-4)(x-1)(x-3)(x-2)+1}\] \[=\sqrt{(x^2-5x+4)(x^2-5x+6)+1}\] \[=\sqrt{[x^2-5x+4][(x^2-5x+4)+(2)]+1}\] \[=\sqrt{(x^2-5x+4)^2+2(x^2-5x+4)+1}\] what's inside the radical sign is a perfect square of the form u^2 + 2u +1 \[=\sqrt{[(x^2-5x+4)+(1)]^2}\] \[=x^2-5x+5\] change x back to 101 and simplify. :)

OpenStudy (anonymous):

Part II, using algebraic concept, \[101^2=(100+1)^2=100^2+2(100)(1)+1^2=10000+200+1=10201\]

OpenStudy (anonymous):

hahaha

OpenStudy (anonymous):

nice trick... but why do u let x=101, may i take x=97,98,99,or 100 ? it will work too?

OpenStudy (anonymous):

I dunno if it will work for other numbers. I encountered a similar problem and what I used was x=101 haha

OpenStudy (anonymous):

okokokok...^^ thanks